Ray and Gary are the same age. The sum of Ray, Gary and Steve's age is 48. Steve is 1/3 the age of Ray and Gary combined. How old is Steve?
Answer by stanbon(75887)

Ray and Gary are the same age. The sum of Ray, Gary and Steve's age is 48. Steve is 1/3 the age of Ray and Gary combined. How old is Steve?
---
Equations:
r = g
-
r + g + s = 48
---
s = (1/3)(r+g)
------
==================
Rearrange and combine to get
(r+g) = 3s
---
3s + s = 48
4s = 48
Steve = 12 yrs. old
====================
Cheers,
Stan H.